# Soban Nasir Lone

### v1

@@ 5 Discussion
While the $\alpha$-VI DeepONet offers significant flexibility, it is important to acknowledge the associated computational overhead. Compared to the deterministic DeepONet, the $\alpha$-VI DeepONet incurs notably higher training times. This increase arises from two primary factors: the approximation of the Rényi $\alpha$-divergence, which necessitates Monte Carlo sampling, and the doubling of trainable parameters due to the estimation of both mean and variance for each weight. However, compared to a standard KLD-VI DeepONet, the training times are approximately 1.1x longer, which is reasonable for the gain in flexibility.
6 Conclusions
@@
In this work, we have introduced a novel, prior-robust, uncertainty-aware DeepONet framework grounded in generalised variational inference (GVI). This framework enables the learning of complex nonlinear operators while providing robust uncertainty quantification. Unlike previous approaches, our framework adopts an optimisation-centric perspective on Bayesian modeling by minimising the GVI objective. By replacing the Kullback-Liebler divergence with Rényi’s $\alpha$ divergence, we enhance the model’s robustness to prior misspecification and achieve superior predictive performance compared to the standard KLD-VI approach. The hyperparameter $\alpha$ offers flexibility in controlling the degree of robustness, allowing adaptation to different problem characteristics.
Our numerical investigation across four benchmark problems consistently demonstrates the advantages of the $\alpha$-VI DeepONet over both deterministic and standard variational methods in terms of NMSE and NLL. The optimal $\alpha$ value varies across problems, highlighting the importance of tuning the hyperparameter. While the $\alpha$-VI DeepONet offers significant improvements, it is not without limitations. The introduction of the $\alpha$ hyperparameter increases model complexity and requires additional computational resources due to the approximation of Rényi’s $\alpha$ divergence. Furthermore, the mean-field assumption underlying the variational posterior may not be adequate in all cases. Future research directions include exploring alternative variational families (e.g., normalising flows) to address the limitations of the mean-field approximation and developing more efficient approximation techniques for Rényi’s $\alpha$-divergence. Additionally, investigating the application of the proposed framework to a wider range of complex systems is a promising avenue for future work.
7 Acknowledgements
R. Nayek acknowledges the financial support received from the Science and Engineering Research Board (SERB) via Grant number SRG/2022/001410, and the Ministry of Ports, Shipping and Waterways.

### v2

@@ 5 Discussion
In our current approach, the optimal $\alpha$ is selected through cross-validation. This strategy was chosen deliberately to provide a systematic mapping of the performance landscape, thereby revealing the trade-offs between mass-covering ($\alpha<1$) and mode-seeking ($\alpha>1$) regimes. While this manual exploration is computationally more demanding, it offers valuable interpretability that cannot easily be obtained through direct optimisation alone. For practical deployment, however, the framework is fully compatible with more efficient hyperparameter optimisation methods. To demonstrate this compatibility, we conducted a preliminary experiment using Bayesian optimisation on the antiderivative example. BO explored the parameter space in steps of $0.1$, efficiently pruning less promising regions and converging to an optimal setting of $\alpha=1.2$ (NMSE $9.17\times 10^{-5}$). This closely matches the best grid search result, obtained at $\alpha=1.25$ (NMSE $6.59\times 10^{-5}$), averaged across 10 systematic experiments. This demonstrates that Bayesian optimisation, gradient-based approaches, or meta-learning methods can efficiently navigate the $\alpha$-landscape while significantly reducing computational cost, retaining the benefits of robust prior-aware inference for practical applications.
@@
While the $\alpha$-VI DeepONet offers significant flexibility, it is important to acknowledge the associated computational overhead. Compared to the deterministic DeepONet, the $\alpha$-VI DeepONet incurs notably higher training times. This increase arises from two primary factors: the approximation of the Rényi $\alpha$-divergence, which necessitates Monte Carlo sampling and the doubling of trainable parameters due to the estimation of both mean and variance for each weight. However, compared to a standard KLD-VI DeepONet, the training times are approximately 5% longer, which is reasonable for the gain in flexibility.
Beyond empirical performance, it is worth situating our contribution within the broader discussion on prior misspecification in Bayesian neural networks. Recent work has shown that commonly used priors, such as isotropic Gaussians, can be unintentionally informative and contribute to phenomena like the cold posterior effect [wenzel2020good, fortuin2021bayesian]. At the same time, alternative perspectives argue that expressive priors can be constructed even when starting from fully factorised Gaussian assumptions [pearce2020expressive]. Our work does not aim to design new priors but instead complements these efforts by providing a prior-robust inference framework. By employing Rényi’s $\alpha$-divergence within the GVI formalism, we mitigate the adverse effects of imperfect priors, thereby offering a practical path forward when richer priors are computationally prohibitive.
In addition to the prior design, the choice of posterior parametrisation plays a critical role in Bayesian DeepONets. More expressive variational families, for example, those based on normalizing flows, or advanced sampling-based techniques, such as replica-exchange SGLD, can, in principle, produce more accurate posterior approximations and potentially improve both predictive accuracy and uncertainty calibration. However, these approaches often come with substantial computational cost, particularly in the high-dimensional setting of operator learning [lin2021accelerated]. In this work, we chose to retain the efficiency and scalability of standard variational inference, focusing instead on improving robustness to prior misspecification through a divergence-based modification. Exploring richer posterior parametrisations in combination with $\alpha$-VI DeepONets is therefore a natural and promising avenue for future work, offering the potential to combine improved accuracy with principled uncertainty quantification at scale.
Another important factor influencing the performance of neural operators is the sampling strategy used to construct the training datasets. In this work, we followed the standard practice of using random uniform sampling, consistent with prior DeepONet and neural operator literature [lin2021accelerated, garg2023vb], to ensure comparability with earlier studies. However, this approach does not explicitly account for regions in the domain where physical constraints or sensitivity to initial/boundary conditions play a disproportionate role in solution accuracy. Incorporating non-uniform or adaptive sampling schemes that allocate more samples to such critical regions could reduce localised errors and improve both predictive accuracy and uncertainty calibration. Investigating these strategies, potentially in conjunction with richer posterior approximation techniques, represents another promising avenue for future research.
6 Conclusion
@@
This work introduces a prior-robust, uncertainty-aware DeepONet framework grounded in generalized variational inference. The proposed approach enables the learning of complex non-linear operators while providing robust uncertainty quantification. Unlike previous methods, our framework adopts an optimisation-centric perspective on Bayesian modelling by minimising the GVI objective. By replacing the Kullback–Leibler divergence with Rényi’s $\alpha$-divergence, we enhance robustness to prior misspecification and achieve improved predictive performance compared to the standard KLD–VI approach. The hyperparameter $\alpha$ further provides a flexible mechanism for controlling the trade-off between robustness and concentration, allowing adaptation to different problem characteristics.
Our numerical investigations across four benchmark problems, supplemented with additional noisy data and out-of-distribution tests, consistently demonstrate the advantages of the $\alpha$–VI DeepONet over both deterministic and standard variational methods in terms of NMSE and NLL. The variation in optimal $\alpha$ values across problems highlights the importance of tuning this hyperparameter and motivates the development of more efficient selection strategies. While our experiments focused on data generated from 1D Gaussian Random Fields, the regression-based nature of our framework makes it directly compatible with higher-dimensional GRFs, including the 2D setting.
While the $\alpha$–VI DeepONet offers clear benefits, it is not without limitations. The introduction of the $\alpha$ hyperparameter slightly increases model complexity and computational cost due to the approximation of Rényi’s $\alpha$-divergence. Moreover, the mean-field assumption underlying the variational posterior neglects potential correlations between network parameters and may limit the expressiveness of the posterior in high-dimensional settings.
Future work should therefore explore more expressive variational families (e.g., normalising flows) or sampling-based alternatives to address the limitations of the mean-field approximation, as well as develop more efficient approximation techniques for Rényi’s $\alpha$-divergence. Extending the application of the proposed framework to a wider range of complex systems is another promising avenue.
Finally, although our experiments focused on DeepONets, the proposed $\alpha$–VI formulation is not tied to this architecture in particular. Because it modifies only the divergence measure in the variational inference step, it can be readily extended to other operator-learning frameworks such as Fourier Neural Operators, physics-informed neural operators, or multiphysics extensions. This generality reinforces the potential of the method as a broadly applicable tool for prior-robust operator learning.
CRediT authorship contribution statement
Soban Nasir Lone: Conceptualization, Methodology, Software, Validation, Visualization, Writing - original draft. Subhayan De: Visualization, Writing - review & editing. Rajdip Nayek: Conceptualization, Supervision, Methodology, Writing - review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The data will be made available upon reasonable request.